# Visibly Fair Mechanisms

### v1

@@ 2 Model and Definitions
    a finite set of officers $I=\{i_{1},i_{2},\ldots\}$,
@@
    a finite set of states $S=\{s_{1},s_{2},\ldots\}$,
@@
    a capacity for each state $(q_{s})_{s\in S}$, such that $\sum_{s\in S}q_{s}\geq|I|$,
@@
    a strict and complete preference for each officer $(\succ_{i})_{i\in I}$ over states $S$, and
@@
For a given problem, the goal is to produce an allocation of officers to states. Formally, an allocation $a=(a_{i})_{i\in I}$ is a list specifying a state $a_{i}\in S$ for each officer $i\in I$. An allocation is feasible if, for each $s\in S$, we have |$|\{i\in I:a_{i}=s\}|\leq q_{s}$. We denote the set of all feasible allocations by $A$.
While the allocation decision is typically based on officers’ reported preferences and assigned priorities, in our setup, officers do not necessarily communicate their full preferences directly but instead provide partial preference information from a menu of partial preferences available.
More specifically, for each officer $i\in I$, let $M_{i}$ denote the message space for $i$. For every message $m_{i}\in M_{i}$, let $\succ_{m_{i}}$ denote the strict partial order it reports:
$s\succ_{m_{i}}s^{\prime}\quad\Longleftrightarrow\quad(s,s^{\prime})\in m_{i}.$
Throughout, “$\succ_{m_{i}}$” will be read as “preferred under the message $m_{i}$.” Each message $m_{i}\in M_{i}$ consists of a strict partial order over the set of states $S$; that is, $m_{i}\subseteq S\times S$ is [^5] A binary relation $R$ on $S$ is a strict partial order if it is irreflexive ($\lnot(sRs)$ for every $s$) and transitive ($sRs^{\prime}\land s^{\prime}Rs^{\prime\prime}\Rightarrow sRs^{\prime\prime}$). irreflexive and transitive. We say that states $s$ and $s^{\prime}$ are comparable under $M_{i}$ if, for every $m_{i}\in M_{i}$, either $s\succ_{m_{i}}s^{\prime}$ or $s^{\prime}\succ_{m_{i}}s$. Moreover, for any $m_{i}\in M_{i}$, we denote by $\succsim_{m_{i}}$ the associated weak preference—that is, $s\succsim_{m_{i}}s^{\prime}\iff s\succ_{m_{i}}s^{\prime}\text{ or }s=s^{\prime}$.
@@
A message space $M_{i}$ satisfies richness if, whenever there exists a message $m_{i}\in M_{i}$ and two states $s,s^{\prime}\in S$ such that $s\,\succ_{m_{i}}\,s^{\prime}$, there also exists a message $m_{i}^{\prime}\in M_{i}$ such that:
@@
In words, whenever the designer allows an officer to send a message that ranks $s$ above $s^{\prime}$, there also exists a message $m_{i}^{\prime}\in M_{i}$ that preserves every other pairwise comparison of $m_{i}$, but reverses the comparison of $s$ and $s^{\prime}$.
Henceforth, we assume each message space $M_{i}$ satisfies richness, unless noted otherwise. This property guarantees that the preference component of the message space is balanced: it never forces the officer to reveal a comparison in one direction without permitting the symmetric comparison in the opposite direction. Whenever the message space is clear from context we continue to write simply $M=(M_{i})_{i\in I}$ and $m=(m_{i})_{i\in I}$.
@@ 3.1 Visible Fairness
    there is a $s\in S$ such that $a_{i}\not=s$, |$|\{i\in I:a_{i}=s\}|<q_{s}$, and $s\succ_{m_{i}}a_{i}$.
@@
At first sight, visible (un)fairness appears to be a combination of standard non-wastefulness and elimination of justified envy. The distinction, however, lies on the fact that $m_{i}$, in general is incomplete, and therefore some existing wastefulness or justified envy is “invisible” due to the limits to the expression of the associated prefrences imposed by the message space.
@@
We next introduce a new family of mechanisms.
@@
A mechanism $\psi^{Q}$ is a $m$-queue allocation mechanism if $\psi^{Q}(m)$ is the outcome produced by the following procedure:
  • Step 00: Set $S^{1}=S$.
  • Step $k$ ($1\leq k\leq n$): $a_{k}=s^{k}\in G(S^{k},m_{k})$. If the number of officers assigned to $s^{k}$ reaches $q_{s^{k}}$, that is |$\left|\{i\in I:i\leq k\text{ and }a_{i}=s^{k}\}\right|=q_{s^{k}}$, then $S^{k+1}\equiv S^{k}\backslash\{s^{k}\}$. Otherwise, $S^{k+1}=S^{k}$.
The $m$-queue allocation mechanism is the natural partial-order analogue of serial dictatorship. Starting from the highest-ranked officer, these are matched to undominated states, given previously assigned ones. The difference here being, of course, that there might be multiple undominated states, and therefore some selection criterion between them must also be defined.
@@
$\psi$ is visibly fair if and only if it is a $m$-queue allocation mechanism.
@@
If for every $M_{i}$ and every $s,s^{\prime}\in S$, states $s$ and $s^{\prime}$ are comparable, then Serial Dictatorship is the unique visibly fair $M$-mechanism.
Theorem 1 provides a full characterization of visibly fair mechanisms: any such mechanism must be an $m$-queue allocation mechanism. While this definition is compact, it captures an extensive class of mechanisms. Designing an $m$-queue mechanism involves two layers of choice—first, specifying the message spaces officers can use, and second, given a message space, determining which of the undominated ($m_{i}$-maximal) states each officer is matched to. This latter choice can depend on exogenous policy parameters or on the overall message profile. In contrast, as indicated in the corollary above, when officers are allowed to submit complete preferences over all states, visible fairness alone pins down a unique mechanism—serial dictatorship. The characterization in Theorem 1 therefore illustrates how relaxing the message space dramatically expands the set of mechanisms that satisfy visible fairness, enabling greater flexibility in accommodating policy goals beyond those implementable via serial dictatorship.
@@ 3.2 Two Special Message Spaces
We will now consider two natural families of message spaces, which are used in some of the existing real-life applications of our concepts and that will serve as the base for the implementation of distributional objectives in Section 5.
@@ 3.2.1 Zonal Message Space
We now introduce a particular family of message spaces that induce a partition of the set of states into “zones.” Formally, suppose a message space $M_{i}$ is such that the states $S$ can be partitioned into disjoint subsets $Z=\{z_{1},\ldots,z_{\ell}\}$ with $\bigcup_{j=1}^{\ell}z_{j}=S$ and $z_{j}\cap z_{j^{\prime}}=\varnothing$ for $j\neq j^{\prime}$. We call $M_{i}$ a zonal message space if, within each subset $z_{j}$, any two states are comparable, while any two states in different subsets are never comparable. Concretely:
Within-zone completeness: For every zone $z_{j}$ and every pair of distinct states $(s,s^{\prime})$ in $z_{j}$, there exists some message $m_{i}\in M_{i}$ such that $s\succ_{m_{i}}s^{\prime}$, and also another message $m_{i}^{\prime}\in M_{i}$ such that $s^{\prime}\succ_{m_{i}^{\prime}}s$. In other words, within each zone, the message space is rich enough to permit any ranking of states within that zone.
@@
The zonal structure will be convenient for discussing distributional objectives in later sections (Section 5). Intuitively, each zone can be treated as an internally ranked “block” of states, allowing certain policy objectives—such as balancing officer assignments across different geographical or administrative regions—to be embedded in how officers are permitted to convey their preferences.
Notice that given any $Z$, $i$, zonal message space $M_{i}$, and a preference over states $\succ_{i}$, there is exactly one message $m_{i}\in M_{i}$ such that for every pair of states $s,s^{\prime}$, $s\succ_{i}s^{\prime}\iff s\succ_{m_{i}}s^{\prime}$. We will say that $m_{i}$ represents $\succ_{i}$ in $M_{i}$.
Suppose each officer has a zonal message space. Consider a zone selection function $\mathcal{C}_{i}:2^{S}\times{M}\to Z$, which is a function such that for any $X\subseteq S$, $X\,\cap\,\mathcal{C}_{i}(X,m)\not=\varnothing$. [^6] That is, as long as there are states with spare capacity available, $C_{i}$ must choose a zone with at least one of them.
@@
  • Step 00: Set $S^{1}=S$.
  • Step $k$ ($1\leq k\leq n$): $a_{k}\in G(S^{k},m_{k})\cap\mathcal{C}_{k}(S^{k},m))=\{s^{k}\}$. [^7] It is easy to see that within each zone, there is a unique $m_{i}$-maximal element, since all states in a zone are comparable. If the number of officers assigned to $s^{k}$ reaches $q_{s^{k}}$, that is |$\left|\{i\in I:i\leq k\text{ and }a_{i}=s^{k}\}\right|=q_{s^{k}}$, then $S^{k+1}\equiv S^{k}\backslash\{s^{k}\}$. Otherwise, $S^{k+1}=S^{k}$.
@@ 3.2.2 Zonal Message Space with Ranking over the Zones
We now enrich the idea of a zonal message space by allowing officers to impose a strict ordering across the zones. As before, let $Z=\{z_{1},\dots,z_{\ell}\}$ be a partition of $S$ into disjoint zones. Let, moreover, $\max(X,m_{i})\equiv\bigl{\{}\,s\in X:\nexists s^{\prime}\in X,\;s^{\prime}\succ_{m_{i}}s\,\bigr{\}}$ and $\min(X,m_{i})\equiv\bigl{\{}\,s\in X:\nexists s^{\prime}\in X,\;s\succ_{m_{i}}s^{\prime}\bigr{\}}$. A zonal message space with ranking over the zones is a message space $M_{i}$ in which:
Within each zone, every pair of states is fully comparable (i.e., for any two states $s,s^{\prime}\in z$, there exist messages in which $s\succ_{m_{i}}s^{\prime}$ and others in which $s^{\prime}\succ_{m_{i}}s$), just as in the standard zonal case.
Across zones, each message $m_{i}\in M_{i}$ implies a ranking over zones:
$z_{1}\;\triangleright_{m_{i}}\;z_{2}\;\triangleright_{m_{i}}\;\dots\;\triangleright_{m_{i}}\;z_{\ell},$
where
@@
In words, a these rankings augment zonal message spaces in a minimal sense: it allows for officers to express some information about how they rank zones in the weakest sense: by ranking zone $z$ above $z^{\prime}$, they are saying that the best state in $z$ is preferred over the worst in $z^{\prime}$. Notice that if this was not the case, we would have $z$ being ranked above $z^{\prime}$ while every state in $z^{\prime}$ is preferred to every state in $z$ [^8] To see this, suppose that $Z=\{z_{1},z_{2}\}$, $z_{1}=\{s_{1},s_{2}\}$ and $z_{2}=\{s_{3},s_{4}\}$. Let $m_{i}$ be such that $s_{1}\succ_{m_{i}}s_{2}$ and $s_{3}\succ_{m_{i}}s_{4}$. If a ranking $z_{1}\;\triangleright_{m_{i}}\;z_{2}$ is associated with $s_{1}\succ_{m_{i}}s_{4}$, we have this additional comparison between states and nothing else. But if it was associated with $s_{4}\succ_{m_{i}}s_{1}$, this would imply that $s_{3}\succ_{m_{i}}s_{4}\succ_{m_{i}}s_{1}\succ_{m_{i}}s_{2}$, that is, a complete preference in which every state in $z_{2}$ is preferred to every state in $z_{1}$, which would be at odds with any reasonable interpretation of what $z_{1}\;\triangleright_{m_{i}}\;z_{2}$ implies for preferences among these states.
@@
Officers would be required to provide complete rankings among $\{s_{1},s_{2}\}$ and among $\{s_{3},s_{4}\}$, and a ranking over zones $\triangleright_{m_{i}}$. Consequently, any message in this space is of the form:
@@
A ranked zone selection function restricts which zone will be assigned to an officer on the basis of the limits visible fairness imply given the ranking over zones: if a zone $z$ contains only the state deemed as the least preferred in $z$, no zone $z^{\prime}$ for which $z^{\prime}\triangleright_{m_{i}}z$ can have capacity in its most-preferred state.
@@
  • Step 00: Set $S^{1}=S$.
  • Step $k$ ($1\leq k\leq n$): $a_{k}\in G(S^{k},m_{k})\cap\mathcal{C}_{k}(S^{k},m))=\{s^{k}\}$. [^9] Ranked zone selection function makes sure that within the selected zone, $m_{k}$-maximal element exists. Since all states in a zone are comparable, this element is unique. If the number of officers assigned to $s^{k}$ reaches $q_{s^{k}}$, that is |$\left|\{i\in I:i\leq k\text{ and }a_{i}=s^{k}\}\right|=q_{s^{k}}$, then $S^{k+1}\equiv S^{k}\backslash\{s^{k}\}$. Otherwise, $S^{k+1}=S^{k}$.

### v2

@@ 2 Model and Definitions
    a finite set of officers $I=\{i_{1},i_{2},\ldots,i_{n}\}$,
@@
    a finite set of states $S=\{s_{1},s_{2},\ldots,s_{m}\}$,
@@
    a capacity for each state $(q_{s})_{s\in S}$, such that $\sum_{s\in S}q_{s}\geq n$,
@@
    a strict preference (asymmetric, complete, and transitive) for each officer $(\succ_{i})_{i\in I}$ over states SS, [^6] We denote by $\succsim_{i}$ the associated weak preference—that is, $s\succsim_{i}s^{\prime}\iff s\succ_{i}s^{\prime}\text{ or }s=s^{\prime}$. and
@@
For a given problem, the goal is to produce an allocation of officers to states. Formally, an allocation $a=(a_{i})_{i\in I}$ is a list specifying a state $a_{i}\in S$ for each officer $i\in I$. An allocation is feasible if, for each $s\in S$, we have $|\{i\in I:a_{i}=s\}|\leq q_{s}$. We denote the set of all feasible allocations by $\mathcal{A}$. Furthermore, without loss of generality, we assume that officers with lower subscripts have a higher priority, i.e., $\pi(i_{1})<\pi(i_{2})<\cdots<\pi(i_{n})$. While the allocation decision is based on officers’ reported preferences and assigned priorities, in our setup, officers do not necessarily communicate their full preferences directly but instead provide partial preference information from a menu of partial preferences available.
[…]
@@
A message space $M_{i}$ satisfies richness if, whenever there exists a message $m_{i}\in M_{i}$ and two states $s,s^{\prime}\in S$ such that $s\,\succ_{m_{i}}\,s^{\prime}$ and there is no state $y\in S$ where $s\,\succ_{m_{i}}\,y\,\succ_{m_{i}}\,s^{\prime}$, then there also exists a message $m_{i}^{\prime}\in M_{i}$ such that:
@@
In words, whenever the designer allows an officer to send a message that ranks $s$ above $s^{\prime}$, there also exists a message $m_{i}^{\prime}\in M_{i}$ that preserves every other pairwise comparison of $m_{i}$, but reverses the comparison of $s$ and $s^{\prime}$. The condition involving the third state $y$ prevents this reversal from violating transitivity.
Henceforth, we assume each message space $M_{i}$ satisfies richness, unless noted otherwise. This property guarantees that the message space never forces the officer to reveal a comparison in one direction without permitting the symmetric comparison in the opposite direction.
@@ 3.1 Visible Fairness
    there is a $s\in S$ such that $a_{i}\not=s$, $|\{i\in I:a_{i}=s\}|<q_{s}$, and $s\succ_{m_{i}}a_{i}$.
@@
At first sight, visible (un)fairness appears to be a combination of standard non-wastefulness and elimination of justified envy. The distinction, however, lies in the fact that $m_{i}$, in general, is incomplete, and therefore some existing wastefulness or justified envy is “invisible” due to the limits to the expression of the associated prefrences imposed by the message space.
@@
Since $\succ_{m_{i}}$ is acyclic and $S$ is finite, $G(X,m_{i})\not$$\varnothing$. [^8] For a proof, see Bossert and Suzumura (2010, Theorem 2.6). We next introduce a new family of mechanisms.
@@
A mechanism $\psi$ is a $\mathbf{m}$-queue allocation mechanism if $\psi(m)$ is the outcome produced by the following procedure:
  • Step 0: Set $S^{1}=S$.
  • Step $k$ ($1\leq k\leq n$): $a_{k}=s^{k}\in G(S^{k},m_{k})$. If the number of officers assigned to $s^{k}$ reaches $q_{s^{k}}$, that is $\left|\{i\in I:i\leq k\text{ and }a_{i}=s^{k}\}\right|=q_{s^{k}}$, then $S^{k+1}\equiv S^{k}\backslash\{s^{k}\}$. Otherwise, $S^{k+1}=S^{k}$.
The $m$-queue allocation mechanism is the natural partial-preference analogue of serial dictatorship. Starting from the highest-ranked officer, they are matched to undominated states, given previously assigned ones. The difference here being, of course, that there might be multiple undominated states, and therefore some selection criterion between them must also be defined.
@@
A mechanism $\psi$ is visibly fair if and only if it is a $m$-queue allocation mechanism.
@@
If for every $i\in I$, every $m_{i}\in M_{i}$ and every $s,s^{\prime}\in S$, states $s$ and $s^{\prime}$ are comparable under $m_{i}$, then Serial Dictatorship is the unique visibly fair $M$-mechanism.
Theorem 1 provides a full characterization of visibly fair mechanisms: any such mechanism must be an $m$-queue allocation mechanism. While this definition is compact, it captures an extensive class of mechanisms. Designing an $m$-queue mechanism involves two layers of choice—first, specifying the message spaces officers can use, and second, given a message space, determining which of the undominated ($m_{i}$-maximal) states each officer is matched to. This latter choice can depend on exogenous policy parameters or on the overall message profile. In contrast, as indicated in the corollary above, when officers are allowed to submit complete preferences over all states, visible fairness alone pins down a unique mechanism—serial dictatorship. The characterization in Theorem 1 therefore illustrates how relaxing the message space, dramatically expands the set of mechanisms that satisfy visible fairness, enabling greater flexibility in accommodating policy goals beyond those implementable via serial dictatorship.
@@ 3.2 Two Special Message Spaces
We now turn to two natural families of message spaces, which are used in practice and serve as the foundation for implementing distributional objectives in Section 5.
@@ 3.2.1 Zonal Message Space
We now introduce a family of message spaces that induce a partition of the set of states into “zones.” Formally, suppose a message space $M_{i}$ is such that the states $S$ can be partitioned into disjoint subsets $Z=\{z_{1},\ldots,z_{\ell}\}$ with $\bigcup_{j=1}^{\ell}z_{j}=S$ and $z_{j}\cap z_{j^{\prime}}=\varnothing$ for $j\neq j^{\prime}$. We call $M_{i}$ a zonal message space if within each subset $z_{j}$, any two states are comparable under any $m_{i}\in M_{i}$, while any two states in different subsets are never comparable under any $m_{i}\in M_{i}$. That is, zonal message spaces have the following properties:
Within-zone completeness: For every strict total order $R$ on $S$, there exists a message $m_{i}\in M_{i}$ such that, for every zone $z_{j}\in Z$ and all states $s,s^{\prime}\in z_{j}$ with $s\neq s^{\prime}$,
$s\,R\,s^{\prime}\quad\Longleftrightarrow\quad s\succ_{m}s^{\prime}.$
In other words, within each zone, the message space is rich enough to permit any ranking of states within that zone.
@@
Notice that given any $Z$, $i$, zonal message space $M_{i}$, and a preference over states $\succ_{i}$, there is exactly one message $m_{i}\in M_{i}$ such that for every pair of states $s,s^{\prime}$, $s\succ_{i}s^{\prime}\iff s\succ_{m_{i}}s^{\prime}$.
Suppose each officer has a zonal message space. Consider a zone selection function $\mathcal{C}_{i}:2^{S}\times{M}\to Z$, which is a function such that for any $X\subseteq S$, $X\,\cap\,\mathcal{C}_{i}(X,m)\not$$\varnothing$. [^9] That is, as long as there are states with spare capacity available, $\mathcal{C}_{i}$ must choose a zone with at least one of them.
@@
  • Step 0: Set $S^{1}=S$.
  • Step $k$ ($1\leq k\leq n$): $a_{k}\in G(S^{k},m_{k})\cap\mathcal{C}_{k}(S^{k},m))=\{s^{k}\}$. [^10] It is easy to see that within each zone, there is a unique $m_{i}$-maximal element among any set of remaining states with spare capacity, since all states in a zone are comparable under $m_{i}$. If the number of officers assigned to $s^{k}$ reaches $q_{s^{k}}$, that is $\left|\{i\in I:i\leq k\text{ and }a_{i}=s^{k}\}\right|=q_{s^{k}}$, then $S^{k+1}\equiv S^{k}\backslash\{s^{k}\}$. Otherwise, $S^{k+1}=S^{k}$.
Zone selection functions constitute the essential component of the definition that results in the large variety of these mechanisms. They indicate, for each profile of messages, which zone will be used to determine an officer’s outcome. This zone can depend on some exogenous parameter, on the allocations of higher-ranked officers and/or the preferences stated by other officers, as well as her own message. Once a zone is determined, however, the state that will be matched to the officer depends only on her preferences between the remaining states in that zone. [^11] Notice, moreover, that the definition of the mechanism requires that $\mathcal{C}_{k}$ chooses a zone with states with spare capacity, which by assumption always exists.
@@ 3.2.2 Zonal Message Space with Ranking over the Zones
We now enrich the idea of a zonal message space by allowing officers to impose a strict ordering across the zones. As before, let $Z=\{z_{1},\dots,z_{\ell}\}$ be a partition of $S$ into disjoint zones. Let, moreover, $\max(X,m_{i})\equiv G(X,m_{i})$ and $\min(X,m_{i})\equiv\bigl{\{}\,s\in X:\nexists s^{\prime}\in X,\;s\succ_{m_{i}}s^{\prime}\bigr{\}}$. A zonal message space with ranking over the zones is a message space $M_{i}$ in which:
Within each zone, states are comparable just as in the standard zonal case. That is, for every strict total order $R$ on $S$, there exists a message $m_{i}\in M_{i}$ such that, for every zone $z_{j}\in Z$ and all states $s,s^{\prime}\in z_{j}$ with $s\neq s^{\prime}$,
$s\,R\,s^{\prime}\quad\Longleftrightarrow\quad s\succ_{m}s^{\prime}.$
Across zones, each message $m_{i}\in M_{i}$ implies a complete ranking $\triangleright_{m_{i}}$ over zones $Z$ s.t.
@@
In words, these rankings augment zonal message spaces in a minimal sense: they allow officers to express some information about how they rank zones in a weakest sense: by ranking zone $z$ above $z^{\prime}$, they are saying that the best state in $z$ is preferred over the worst in $z^{\prime}$. Notice that if this was not the case, we would have $z$ being ranked above $z^{\prime}$ while every state in $z^{\prime}$ is preferred to every state in zz. [^12] To see this, suppose that $Z=\{z_{1},z_{2}\}$, $z_{1}=\{s_{1},s_{2}\}$ and $z_{2}=\{s_{3},s_{4}\}$. Let $m_{i}$ be such that $s_{1}\succ_{m_{i}}s_{2}$ and $s_{3}\succ_{m_{i}}s_{4}$. If a ranking $z_{1}\;\triangleright_{m_{i}}\;z_{2}$ is associated with $s_{1}\succ_{m_{i}}s_{4}$, we have this additional comparison between states and nothing else. But if it was associated with $s_{4}\succ_{m_{i}}s_{1}$, this would imply that $s_{3}\succ_{m_{i}}s_{4},s_{4}\succ_{m_{i}}s_{1},s_{1}\succ_{m_{i}}s_{2}$, thus making $s_{3}$ the $m_{i}$-maximal element in $\{s_{1},s_{2},s_{3},s_{4}\}$, which would be at odds with any reasonable interpretation of what $z_{1}\;\triangleright_{m_{i}}\;z_{2}$ implies for preferences among these states.
@@
Officers’ messages contain complete rankings among $\{s_{1},s_{2}\}$ and among $\{s_{3},s_{4}\}$, and a ranking over zones $\triangleright_{m_{i}}$. Consequently, any message in this space is of the form:
@@
A ranked zone selection function restricts which zone will be assigned to an officer on the basis of the limits visible fairness imply given the ranking over zones: if a zone $z$ contains only the state deemed as the least preferred in $z$, no zone $z^{\prime}$ for which $z^{\prime}\triangleright_{m_{i}}z$ can have spare capacity in its most-preferred state.
@@
  • Step 0: Set $S^{1}=S$.
  • Step $k$ ($1\leq k\leq n$): $a_{k}\in G(S^{k},m_{k})\cap\mathcal{C}_{k}(S^{k},m))=\{s^{k}\}$. [^13] Ranked zone selection function makes sure that within the selected zone, $m_{k}$-maximal element exists. To see why, let $z=C_{k}(S^{k},m)$ be the selected zone. We have two cases from the ranked zone selection function definition: Case 1: $S^{k}\cap z\neq\varnothing$ and $S^{k}\cap z\neq\{\min(z,m_{k})\}$ Then $S^{k}\cap z$ contains non-minimal elements of zone $z$. Since all states in a zone are comparable under $m_{k}$, the $m_{k}$-maximal state within $S^{k}\cap z$ is unique. By the across-zone ranking property, only $\min(z,m_{k})$ can be dominated by states in different zones, so the $m_{k}$-maximal state within $S^{k}\cap z$ is also $m_{k}$-maximal in all of $S^{k}$. Therefore, $G(S^{k},m_{k})\cap\mathcal{C}_{k}(S^{k},m))$ is non-empty and a singleton. Case 2: $S^{k}\cap z\neq\varnothing$ and no zone $z^{\prime}\triangleright_{m_{i}}z$ has $\max(z^{\prime},m_{k})\in S^{k}$ The condition ensures that all maximal elements of higher-ranked zones are unavailable in $S^{k}$. Therefore, $m_{k}$-maximal state within $S^{k}\cap z$ faces no domination from higher-ranked zones and is $m_{k}$-maximal in $S^{k}$. Therefore, again $G(S^{k},m_{k})\cap\mathcal{C}_{k}(S^{k},m))$ is non-empty and a singleton. If the number of officers assigned to $s^{k}$ reaches $q_{s^{k}}$, that is $\left|\{i\in I:i\leq k\text{ and }a_{i}=s^{k}\}\right|=q_{s^{k}}$, then $S^{k+1}\equiv S^{k}\backslash\{s^{k}\}$. Otherwise, $S^{k+1}=S^{k}$.
@@
The presence of ranking over zones implies some restrictions on the zones that the zone selection function can determine, as shown in the example below.
Example 3.3.
There are three states $S=\{s_{1},s_{2},s_{3}\}$ (capacity 1 each) and two zones
$z_{1}=\{s_{1}\},\qquad z_{2}=\{s_{2},s_{3}\}.$
There are two officers, $i_{1}$ and $i_{2}$. We will consider visibly fair mechanisms in which the message space is zonal with rankings over these zones for both officers.
Officers $i_{1}$ and $i_{2}$ both submit the same message:
$z_{1}\;\triangleright_{m_{i}}\;z_{2},\qquad s_{2}\succ_{m_{i}}s_{3}.$
Because $z_{2}$ contains two available states, the ranked‑zone selection function could, in this scenario, place $i_{1}$ in either zone $z_{1}$ or $z_{2}$. Suppose that it places on $z_{2}$. Given $m_{1}$, $i_{1}$ is matched to $s_{2}$.
Now $z_{1}$ still has $s_{1}$ free, and $z_{2}$ only $s_{3}$. Since $z_{1}$ remains vacant and is the highest‑ranked for $i_{2}$, the mechanism must choose $z_{1}$ and assign $s_{1}$. Selecting $z_{2}$ would contradict $z_{1}\triangleright_{m_{2}}z_{2}$ while $z_{1}$ still offers an available seat, and is therefore not allowed.
Ordinary partitioned priority mechanisms (without the zone ranking) would leave the planner free to swap $i_{2}$ between $z_{1}$ and $z_{2}$, illustrating how adding cross‑zone orderings tightens the designer’s hands, in comparison.